# Global Prediction of COVID-19 Variant Emergence Using Dynamics-Informed Graph Neural Networks

### v2

@@ IV-D Benchmark Models
Mean Model – Delay Prediction: This model predicts delays based on average delays on the variants between the countries under consideration. Specifically, at time $t$, for a new variant $i$ on a target country $p$, suppose $Q$ is the set of all countries where the variant has appeared in. Then, the predicted time of arrival is given by
${t^{p}_{i}=t+\mbox{median}_{q}\left(\mbox{mean}_{j}\left(\Delta_{j}(p,q)\right)\right)\,.}$ (19)
@@
Here $\Delta_{j}(p,q)$ is the delay between the appearance of variant $j$ from country $q$ to country $p$. We take the mean of $\Delta_{j}(p,q)$ over all prior variants $j$ that appeared in both countries $p$ and $q$. Finally, we take the median of the delays over all countries in $Q$ where variant $i$ has already appeared.
Trivial Model – Emergence Prediction: This model sets all predictions to 1. In other words, this model consistently predicts that a variant will emerge in all countries
@@
Dynamics-based Model – Delay Prediction: This is the dynamics-based linear model discussed in Section III-C. This model serves as a baseline mechanistic model, which, under the right assumptions, offers an explainable and efficient method for calculating delays.
Adjacency-based Model – Emergence Prediction: We use a simple adjacency-based classifier. At a given timestep, $\tau$, and variant, $v$, we check all countries in which $v$ is dominant. If it is dominant in a country $c$, we predict that the variant $v$ will become dominant in all countries adjacent to $c$ at $\tau+1$.
@@
Decision Tree: We use Decision Tree classifiers and regression models to address Emergence Prediction and Delay Prediction, respectively. As discussed in Section IV-B, this involves transforming the dataset into an $N\times 5$ matrix. Here, $N$ represents the total number of samples, where each sample corresponds to the 5 features of a given variant at a specific point in time in a specific country. These trees use the Gini impurity and mean squared error for their criterion functions. No maximum depth was set for either.
@@
Multi Layer Perceptron: We employ a Multi-Layer Perceptron (MLP) classifier and regressor for Emergence Prediction and Delay Prediction, respectively. They were configured with two hidden layers, consisting of 16 and 8 neurons, respectively. The selection of hyperparameters, including the number of hidden layers, activation function, and solver, was determined through comprehensive experimentation and optimization tailored to our specific problem. The training procedure closely followed the approach outlined in Section III-E, with the only distinction being the structuring of our dataset into an $N\times 5$ matrix, as described previously. A batch size equivalent to the number of countries was used, where each batch represented the list of countries and their features at each timestep for each variant, ensuring that the gradient updates were performed collectively for each group.
@@
Gated Recurrent Unit Network We utilize a simple GRU network to encode the temporal information of variant transmissions and predict their emergence and delay. To train this model, first, we feed temporal features of the variant to the GRU architecture which embeds them for each country distinctly. Then, we flatten the embedded features and append them with $S$ and directly give them to an FC layer for predicting the time it takes for that variant to appear in a specific country. We employ GraphNorm layer here after the concatenation to accelerate training by smoothing graph aggregation distributions.
@@
Graph Convolutional Network (GCN) We consider a simple GCN, with a basic architecture in an attempt to capture local relationships and features within the graph structure. It is comprised of two GCN layers, which output 32 and 16 channels respectively. Dropout is strategically employed between these two layers for regularization. The final graph embedding is flattened and fed into an FC layer which outputs the delay or emergence for each country.
@@
Encoded Dynamics GCN Since each country possesses a variable number of circulating variants with different growth rates $S$ (as derived in III-B), we can learn a latent vector representation $Z$ of this data through the use of encoders. We achieve this in two ways:
Using an Autoencoder (AE-GCN): We employ a simple autoencoder with two linear layers and ReLU activation in the encoder and decoder blocks.
The model takes a variable-sized input $S_{pt}$, a subset of variants (retrospective) in a region $p$ at a given time instance $t$. We train the autoencoder using MSE loss. We append the learned encoding from this training process to the feature matrix before constructing a spatio-temporal snapshot graph data for the GCN model (defined earlier).
@@
Embedded Encoder (EE-GCN): Using an autoencoder implies that the model is guided towards learning an embedding using the decoder output, and these two are inherently separate models. The encoder weights are also unaffected during graph training. An alternative approach is to embed this into the larger GCN model. In this method, as before, the variable-sized $S_{pt}$ is fed to an encoder block (consisting of two linear layers with ReLU activation). Instead of using a decoder, we directly concatenate the resulting latent vector representation $Z$ to the node features fed into the GCN. This means the encoder weights are updated in the graph training loop, allowing the model to learn a better latent representation.
IV-E Results
IV-E1 Emergence Prediction
TABLE I: Emergence Prediction f1-scores
Model f1
Trivial Model 0.20
Adjacency-based Model 0.58
Decision Tree 0.96
MLP 0.93
GRU 0.96
GCN 0.73
T-GCN 0.77
AE-GCN 0.73
EE-GCN 0.74
Table I shows the macro F1-scores for emergence prediction across different models. Notably, the Decision Tree and GRU models exhibit high F1-scores, outperforming all other models. While the T-GCN achieves a reasonable performance here relative to the adjacency-based and trivial models, it does not come close to competing with the simpler ML-based models or the GRU. This suggests that identifying the emergence of a variant can be more straightforward and often involves distinct, separable patterns that decision trees and GRUs can efficiently learn.
IV-E2 Delay Prediction
TABLE II: Delay Prediction Errors
@@ IV-E2 Delay Prediction
Mean Model 21.18 26.73
Dynamics-based Model 17.25 34.1
Decision Tree 22.91 23.63
MLP 24.98 23.91
GRU 20.85 22.85
GCN 19.09 21.73
T-GCN 18.45 19.08
AE-GCN 19.77 22.77
EE-GCN 21.00 24.46
DIL-GCN 20.33 22.71
Table II presents the performance metrics for delay prediction across the various models. The evaluated metrics include MedMedAE and MedMAE, as discussed in Section IV-B. The results indicate that the dynamics-based model, employing a dynamics-based linear approach, outperforms all other models in terms of MedMedAE, with T-GCN performing the next best. Additionally, the proposed T-GCN achieves the best MedMAE, showcasing its effectiveness in capturing temporal dependencies for delay prediction, while the dynamics-based model performs worse than all other models in this metric. This suggests that the dynamics-based model is susceptible to outlier predictions for some country pairs, which skews its results. Additionally, we observe that the DIL-GCN does not surpass its GCN counterpart, highlighting the unnecessary use of dynamics as a regularization of the loss when employing the correct features.
@@ V Discussion
(a) T-GCN Median Across Countries
(b) Dynamics-based Model Median Across Countries
Figure 6: Results matrices for the best-performing delay prediction models. The heatmaps display both the errors and the number of countries being predicted for at each timestep until they are no longer dominant anywhere. This is important for analysis as variants may not reach all countries.
In this section, we conduct a detailed examination of our results, show the effectiveness of integrating variant dynamics into our models, and provide justification for the construction of our graph. Note that since high performance was achieved for emergence prediction, this analysis will primarily focus on delay prediction.
@@ V-A Results Analysis
Figure 6 provides a detailed representation of the results for the dynamics-based model and T-GCN model. The heatmaps illustrate the number of countries eligible for predicting at each timestamp. The numbers indicate the errors over time for each variant throughout their existence. Some variants exhibit NaN values for the dynamics-based model. This occurs due to two primary reasons. Firstly, the linear model requires a minimum of two common variants between two countries to build an effective model. Secondly, the linear model can encounter significant challenges when variant $j$ is identical in growth rate to variant $k$ (i.e., $S_{j}=S_{k}$). In such cases, the linear model fits a plane solely along $Z_{1}=0$ as seen in Equation 14. When variant $j$ is no longer identical to variant $k$ for another variant, this model fails drastically, as $Z_{1}\neq 0$. These scenarios were more prevalent in the early stages of the pandemic when only a few variants were circulating, leading to their exclusion from the dynamics-based model’s training process. All final errors, for all models in Section IV-D, were calculated excluding these variants to ensure a fair comparison between the models.
The observations from Figure 6(a) give us two key insights. Firstly, the primary source of error is associated with variants that persist for an extended duration, such as 21J.Delta or 23E.Omicron. Examining the color gradients, these variants appear to linger in only a few countries before spreading to others. In reality, variants typically do not endure for such prolonged periods without either being superseded by another variant or transmitting to additional countries [Chen2023, Verywell_Health_2023]. This suggests a potential issue in the data capturing these lingering variants. Secondly, despite this being a major source of error for GCN-based models, they exhibit overall robustness to outlier predictions, unlike the dynamics-based model as seen in Figure 6(b). Also, note that our ground truth is at a granularity of 14 days which may inflate the error. Suppose, in reality, the first day of a variant reaching a desired prevalence threshold is $t=14k+1$, for some integer $k$. In the ground truth, it will appear at the next multiple of 14, i.e., $t+13$. If our prediction was $t^{\prime}=14k=t-1$, the actual error is 1 day, while in comparison to the observed ground truth, it will result in an error of 14 days.
@@
While the dynamics-based model does output very good predictions, it fails often and dramatically in some cases. Figure 7 presents examples of 3D points in space utilized for fitting linear models. In Figure 7(a), the 3D points form a space where a plane can be fitted exceptionally well. However, as newer COVID variants emerge, the dynamics may evolve, causing some initial assumptions, such as full cross-immunity [Chen2023], to become invalid. This evolution makes it more challenging to perfectly fit a linear model, as depicted in Figure 7(c).
V-B Ablation Studies
In this section, we study the effectiveness of our design choices. While MedMedAE was used as our metric of choice, MedMAE also provided similar results here.
@@ V-B1 Dynamics
TABLE III: Dynamics Ablations MedMedAE
@@
Decision Tree 22.91 27.36 23.39 25.77 23.86
MLP 24.98 26.73 24.34 25.93 23.86
GRU 20.85 21.05 21.05 23.27 n/a
GCN 19.09 21.00 20.68 21.00 20.36
T-GCN 18.45 19.73 22.27 20.36 n/a
We examine the usefulness of integrating the extracted dynamics as inputs for our models. In this analysis, we train the same models using different sets of features. Specifically, we investigate the effectiveness of using $S$ as a feature and the impact of using a time series of $r$. The feature configuration $[r|S]$ is detailed in Section III-D1, while $r$ excludes $S$ as a feature, relying solely on the time series of prevalence ratios rather than the prevalence $p$. Additionally, we further investigate the effectiveness of $S$ by comparing $[p|S]$ to $p$ and the utilization of $S$ as the only feature.
From the results in table III, we observe that the $[r|S]$ combination outperformed all other configurations except for the case of the MLP. There is also an improvement when utilizing $r$ over $p$ for the temporal-based models, with the GCN performing similarly and the others exhibiting poorer performance. This suggests that the temporal models effectively encode the latent information contained in the time series of $r$, while the simpler models struggle to do so. Additionally, the inclusion of $S$ as a feature enhances results across all cases. Even when used in isolation, $S$ performs generally better than the models relying solely on a time series. Combining $S$ and $r$ significantly improves performance in most cases, indicating a synergistic effect likely stemming from the linear relationship between the two, as discussed in Section III-D1. It is important to note that the GRU and T-GCN were not trained using just $S$ since excluding the time series encoding reduces these models to a MLP and GCN, respectively.
@@ V-B2 Graphs
TABLE IV: Graph Ablation Results
@@
GCN 19.09 21.73
GCN_30 20.67 23.29
GCN_60 21.64 23.36
GCN_flights_adj 24.82 23.36
GCN_flights 24.18 24.00
GCN_fullcon 26.09 25.29
GCN_WoEW 20.68 23.34
@@
When designing the structure of our graph, several considerations were taken. Firstly, we acknowledged that certain nodes (countries) might be more susceptible to noise due to reporting inconsistencies or a lack of reporting, leading to unreliable data. To address this issue, we explored the option of using only countries with consistently reported data, focusing on the top 30 (GCN_30) and top 60 (GCN_60) countries out of the total 87 (GCN) countries in our dataset. Additionally, we examined different approaches for creating edges in our graph. While country adjacency served as the base (GCN), we also explored the option of connecting countries through flight routes using the OpenFlights dataset. This exploration included a combination of country adjacency and flight routes (GCN_flights_adj), as well as utilizing only flight routes (GCN_flights). We also include a fully connected graph (GCN_fullcon) for comparison. All these graphs incorporated dynamic weights for their edges, with the exception of GCN_WoEW, which was utilized as a comparison. These ablations were conducted using a simple two-layer GCN to isolate and understand the impact of individual components or features on the overall model performance.
From Table IV, it is evident that our selected graph design, based on country adjacency, outperforms alternative configurations. The number of nodes (countries) in the graph does not significantly impact the model performance. In contrast, an FC graph performs the worst, underscoring the importance of considering specific relationships between countries rather than creating direct connections between all nodes. Interestingly, the use of flight routes generally results in increased errors. While this might seem counter-intuitive, considering that flights could connect continents, it’s crucial to note that, based on the definition of country adjacency in Section IV-A, countries like the USA and Russia are considered adjacent. As a result, they act as a link between continents. Additionally, the inclusion of dynamic weights in the edges proves beneficial for the overall model performance.
V-B3 Dominance
TABLE V: Threshold Ablation Results for GCN
$\theta$ MedMedAE MedMAE f1
$\frac{1}{9}$ 33.41 37.81 0.77
$\frac{1}{4}$ 20.68 24.30 0.75
$\frac{1}{3}$ 19.09 21.73 0.73
$\frac{1}{2}$ 18.45 20.88 0.76
1 17.5 21.56 0.73
We established our dominance threshold, denoted as $\theta$, with a value of $\frac{1}{3}$ in Section III-A. The impact of varying $\theta$ is presented in Table V. While this variation does not significantly influence the problem of emergence prediction, it does play a role in delay prediction. We observe that increasing values of $\theta$ corresponds to decreasing errors. This phenomenon is likely attributed to the fact that the arrival of a variant heavily depends on its status in neighboring countries. Once a variant arrives, its potential to achieve dominance or not simply hinges on its behavior in other countries. This explains why emergence prediction is less affected. However, delay prediction is more complex. Following a variant’s arrival in a country, its trajectory to higher prevalence is largely influenced by internal factors such as the mobility within the country and the precautionary measures implemented by the government to impede its spread. A larger threshold provides the model with more time to leverage these internal dynamics, resulting in more accurate predictions for delay.

### v3

@@ IV-E Benchmark Models
Mean Model – Delay Prediction: This model predicts delays based on average delays on the variants between the countries under consideration. Specifically, at time $t$, for a new variant $i$ on a target country $Y$, suppose $Q$ is the set of all countries where the variant has appeared. Then, the predicted time of arrival is given by
${t^{Y}_{i}=t+\mbox{median}_{q}\left(\mbox{mean}_{j}\left(\Delta_{j}(Y,q)\right)\right)\,.}$ (16)
@@
Here $\Delta_{j}(Y,q)$ is the delay between the appearance of variant $j$ from country $q$ to country $Y$. We take the mean of $\Delta_{j}(Y,q)$ over all prior variants $j$ that appeared in both countries $Y$ and $q$. Finally, we take the median of the delays over all countries in $Q$ where variant $i$ has already appeared.
@@
Dynamics-based Model – Delay Prediction: This is the derived dynamics-based linear model discussed in the previous section. This model serves as a baseline mechanistic model, which, under the right assumptions, offers an explainable and efficient method for calculating delays.
@@
Decision Tree: We use Decision Tree regression models to address Delay Prediction. This involves transforming the dataset into an $N\times 5$ matrix. Here, $N$ represents the total number of samples, where each sample corresponds to the 5 features of a given variant at a specific point in time in a specific country. Meaning $N=T\times V\times C$, where $T$ is the total number of timesteps, $V$ is the total number of variants, and $C$ is the total number of countries. These trees use the Gini impurity and mean squared error for their criterion functions. No maximum depth was set for either.
@@
Multi Layer Perceptron: We employ a Multi-Layer Perceptron (MLP) regressor for Delay Prediction. They were configured with two hidden layers, consisting of 16 and 8 neurons, respectively. The selection of hyperparameters, including the number of hidden layers, activation function, and solver, was determined through comprehensive experimentation and optimization tailored to our specific problem. The training procedure closely followed the approach outlined in Section 3.3, with the only distinction being the structuring of our dataset into an $N\times 5$ matrix, as described for the decision tree. A batch size equivalent to the number of countries $C$ was used, where each batch represented the list of countries and their features at each timestep for each variant, ensuring that the gradient updates were performed collectively for each group of variants at a given timestep.
@@
Gated Recurrent Unit Network We utilize a simple GRU network to encode the temporal information of variant transmissions and predict theirdelay. To train this model, first, we feed temporal features of the variant to the GRU architecture which embeds them for each country distinctly. Then, we flatten the embedded features and append them with $S$ and directly give them to an FC layer for predicting the time it takes for that variant to appear in a specific country. We employ GraphNorm layer here after the concatenation to accelerate training by smoothing graph aggregation distributions.
@@
Graph Convolutional Network (GCN) We consider a simple GCN, with a basic architecture in an attempt to capture local relationships and features within the graph structure. It is comprised of two GCN layers, which output 32 and 16 channels respectively. Dropout is strategically employed between these two layers for regularization. The final graph embedding is flattened and fed into an FC layer which outputs the delay for each country.
@@
Encoded Dynamics GCN Each country has a variable number of circulating variants, each with different growth rates $S$ (as derived earlier). These growth rates can be concatenated and represented as a feature vector $S_{qt}$, which correspond to the variants present in a region $q$ at time $t$. We can learn a latent vector representation $Z_{qt}$ of this data through the use of encoders. We implement this in two ways:
Using an Autoencoder (AE-GCN): We employ a simple autoencoder with two linear layers and ReLU activation in the encoder and decoder blocks. The model takes the feature vector $S_{qt}$ as input. We train the autoencoder using MSE loss to reconstruct the feature vector $S_{qt}$. After training, we take the intermediate latent encoding $Z_{qt}$ output from the encoder and append it to the feature matrix before constructing a spatio-temporal snapshot graph data for the GCN model.
@@
Embedded Encoder (EE-GCN): Using an autoencoder implies that the model is guided towards learning an embedding using the decoder output, and these two are inherently separate models. The encoder weights are also unaffected during graph training. An alternative approach is to embed this into the larger GCN model. In this method, as before, the variable-sized $S_{qt}$ is fed to an encoder block (consisting of two linear layers with ReLU activation). Instead of using a decoder, we directly concatenate the resulting latent vector representation $Z_{qt}$ to the node features fed into the GCN. This means the encoder weights are updated in the graph training loop, allowing the model to learn a better latent representation.
IV-F Results
In this section, we present our key results for Emergence Delay Prediction. We previously noted that prior research has concentrated on predicting the prevalence of a variant after its arrival in a specific region, with logistic regression demonstrating effectiveness in estimating its prevalence over time. However, as logistic regression can manage this task relatively easily, our evaluation focused solely on forecasting a variant’s prevalence before its arrival in a specific region. In other words, during evaluation, we excluded cases where a variant is present in a region but has not yet reached $\theta$.
Numerous models were tested and configured as benchmarks to test against the performance of our proposed method. Other implementations include a trivial Mean Model, the baseline Dynamics-based Model, a Decision Tree, a Multilayer Perceptron (MLP), a GRU, and a GCN without an adjusted loss function. In brief, the Mean Model predicts delays based on the average delays of all prior variants between all the countries in which the variant appeared and our intended country. The Decision Tree, MLP, and GRU regression models make use of $T$ extra features which consist of the average $\vec{r}_{c,i}$ of neighboring countries transforming the dataset into a matrix of size $N\times(2T+1)$. Here, $N=D\times V\times C$ is the total number of samples, $D$ is the total number of retrospective dates, $V$ is the total number of variants, and $C$ is the total number of countries. The full details on these models can be found in the Appendix.
To measure the performance of the models, we use the defined Mean/Median Absolute Errors (MAE/MedAE) in the Appendix and measure the performance over time by finding the mean and median of these errors as:
$\text{MedMAE}_{v}=\text{median}(\text{MAE}_{v,t})\mbox{, }$
$\text{MedMedAE}_{v}=\text{median}(\text{MedAE}_{v,t}),$
where $t$ ranges from 1 to $T_{D}$, $T_{D}$ being the total number of timesteps for which variant $v$ circulates before reaching total dominance, i.e. it globally reaches all its targets and begins to disappear. Finally, the mean across all the variants $v$ gets our two error metrics: MedMAE, and MedMedAE.
TABLE I: Prediction Errors
@@ IV-F Results
Mean Model 3.2 3.87
Dynamics-based Model 1.86 2.9
Decision Tree 3.48 3.43
MLP 2.38 2.48
GRU 1.5 1.37
GCN 1.32 1.43
FA-GCN 1.27 1.29
PI-GCN 2.24 2.01
AE-GCN 1.82 2.25
EE-GCN 2.00 2.49
Table I presents the models’ performance metrics for Delay Prediction. The results indicate that our proposed FA-GCN model outperforms all other models in terms of MedMedAE and MedMAE, showcasing its effectiveness in capturing temporal dependencies for Delay Prediction. The dynamics-based model also has a significantly worse MedMAE than MedMedAE. This suggests that the dynamics-based model is susceptible to outlier predictions for some country pairs, which skews its results. Additionally, we observe that the PI-GCN does not surpass its GCN counterpart, highlighting the unnecessary use of dynamics as a regularization of the loss when employing the correct features.
Figure 6: The heatmaps display both the errors and the number of countries being predicted for each variant at each timestep until they are no longer dominant anywhere.
@@ V-A Results Analysis
Figure 6 provides a detailed representation of the results for the baseline dynamics-based model and the FA-GCN model. The heatmaps illustrate the number of countries eligible for prediction at each timestamp. The numbers indicate the errors over time for each variant throughout their existence. Notably, several variants in the dynamics-based model are denoted by *, signifying instances of model or evaluation failure. These failures stem from three primary factors. Firstly, the linear model requires a minimum of two common variants between two countries to build an effective model. Secondly, the linear model can encounter significant challenges when variant $j$ is identical in growth rate to variant $k$ (i.e., $S_{j}=S_{k}$). In such cases, the linear model fits a plane solely along a fixed axis and will thus always predict some constant $\frac{C}{S_{ij}}$, as seen in Equation 11. But as time passes and variant $j$ is no longer identical to variant $k$ for another variant, this model fails drastically since it would still fit on a fixed axis. These scenarios were more prevalent in the early stages of the pandemic when only a few variants were circulating, leading to their exclusion from the dynamics-based model’s training process. To ensure fair comparisons, all final errors across all models were calculated excluding these variants. Finally, there are cases where a variant appears in a country but fails to reach the threshold $\theta$. These instances were disregarded during the evaluation of all models as they were considered trivial.
The observations from Figure 6 give us two insights. Firstly, the primary source of error is associated with variants that persist for an extended duration, such as 21J.Delta or 23E.Omicron. Examining the color gradients, these variants appear to linger in only a few countries before spreading to others. In reality, variants typically do not endure for such prolonged periods without either being superseded by another variant or transmitting to additional countries [30, 38]. This suggests a potential issue in the data capturing these lingering variants. Secondly, despite this being a major source of error for FA-GCN, it exhibits overall robustness to outlier predictions, unlike the dynamics-based model as seen in Figure 6. This, in turn, severely affects the results of the PI-GCN model.
@@
While the dynamics-based model does output good predictions, it fails often and dramatically in some cases. Figure 7 presents examples of 3D points in space utilized for fitting linear models. In Figure 7(a), the 3D points form a space where a plane can be fitted exceptionally well. However, as newer COVID variants emerge, the dynamics may evolve, causing some initial assumptions, such as full cross-immunity [30], to become invalid. This evolution makes it more challenging to perfectly fit a linear model, as depicted in Figure 7(c).
V-B Ablation Study
@@ V-B1 Dynamics
TABLE II: Feature Ablations MedMedAE
@@
Decision Tree 3.48 3.0 2.81 3.0 2.86
MLP 2.38 2.67 2.69 2.55 2.76
GRU 1.50 1.50 1.95 2.32 n/a
GCN 1.32 1.50 1.78 2.32 2.38
FA-GCN 1.27 1.55 1.73 1.73 n/a
We examine the usefulness of integrating the extracted dynamics as inputs for our models. In this analysis, we train the same models using different sets of features. Specifically, we investigate the effectiveness of using $S$ as a feature and the impact of using a time series of $r$. Where $r$ excludes $S$ as a feature, relying solely on the time series of prevalence ratios rather than the prevalence $p$. Additionally, we investigate the effectiveness of $S$ by comparing $[p|S]$ to $p$ and utilizing $S$ as the only feature.
From the results in Table II, we observe that the $[r|S]$ combination outperformed all other configurations except for the case of the Decision Tree. There is also an improvement when utilizing $r$ over $p$ for the temporal-based models, with the GCN performing similarly and the others exhibiting poorer performance. This suggests that the temporal models effectively encode the latent information contained in the time series of $r$, while the simpler models struggle to do so. Additionally, inclusion of $S$ as a feature enhances results across all cases. Combining $S$ and $r$ significantly improves performance in most cases, indicating a synergistic effect likely stemming from the linear relationship between the two, as previously discussed. It is important to note that the GRU and FA-GCN were not trained using just $S$ since excluding the time series encoding reduces these models to an MLP and GCN, respectively.
@@ V-B2 Graphs
TABLE III: Graph Ablation Results
@@
FA-GCN 1.27 1.29
FA-GCN_30 1.46 1.45
FA-GCN_60 1.57 1.46
FA-GCN_flights_adj 1.95 1.46
FA-GCN_flights 1.88 1.53
FA-GCN_fullcon 2.10 1.67
FA-GCN_WoEW 1.46 1.46
@@
When designing the structure of our graph, several considerations were taken. Firstly, we acknowledged that certain nodes (countries) might be more susceptible to noise due to reporting inconsistencies or a lack of reporting, leading to unreliable data. To address this issue, we explored the option of using only countries with consistently reported data, focusing on the top 30 (FA-GCN_30) and top 60 (FA-GCN_60) countries out of the total 87 (FA-GCN) countries in our dataset.
Additionally, we examined different approaches for creating edges in our graph. While country adjacency served as the base (FA-GCN), we also explored the option of connecting countries through flight routes using the OpenFlights dataset. This exploration included a combination of country adjacency and flight routes (FA-GCN_flights_adj), as well as utilizing only flight routes (FA-GCN_flights).
To ensure a thorough comparison, we also tested a fully connected graph (FA-GCN_fullcon), where every country was connected to every other country. This configuration was expected to test the limits of connectivity by removing the specificity of relationships, allowing us to see if more generalized connections might yield better results. Additionally, we included a model without edge weights (FA-GCN_WoEW) to serve as a baseline, assessing the importance of dynamic weights in the performance of the graph model.
From Table III, it is evident that our selected graph design, based on country adjacency, outperforms alternative configurations. The number of nodes (countries) in the graph does not significantly impact the model performance, showcasing its robustness. In contrast, a fully connected graph performs the worst, indicating that simply increasing the number of connections without considering the relevance of those connections can introduce noise and degrade the model’s accuracy. It emphasizes the importance of meaningful relationships between nodes, rather than indiscriminately connecting all nodes.
Interestingly, the use of flight routes generally results in increased errors when compared to the adjacency-based model. While this might seem counter-intuitive, considering that flights could connect distant countries and continents. It’s crucial to note that, based on the definition of country adjacency used by the dataset, countries like the USA and Russia are considered adjacent. As a result, they act as a link between continents. In such cases, the adjacency-based connections might already encapsulate the critical pathways for information flow, making the additional flight route data redundant or even misleading. Moreover, the incorporation of dynamic weights appears to improve model performance, as seen when comparing FA-GCN with FA-GCN_WoEW.
Overall, these results underline the importance of carefully considering the graph structure and the nature of the connections when designing GCN models. While flight routes might seem like a logical addition, their effectiveness depends on how well they complement the existing adjacency information. In this case, it appears that the adjacency-based graph design captures the essential relationships better than the other configurations.